OpenStudy (anonymous):

You and 2 friends visit a pizza shop. There are 12 toppings. 7 toppings are meat and the rest are vegetable. each of you order a different topping (no repeats) at random what is the probability that your group orders only meat

OpenStudy (anonymous):

I like this type of questions because I always get confused so easily. 7 meat toppings and 5 veggie toppings and we pick 3 because of "you and 2 friends." the probability that your group orders only meat is \[\left(\begin{matrix}12 \\ 7\end{matrix}\right)\left(\begin{matrix}7 \\ 3\end{matrix}\right)\] Is that right?